OpenStudy (anonymous):

4 7/8 + 4 5/8 write as a mixed number what is a mixed number? 8 12/8??

OpenStudy (anonymous):

\[4\frac{7}{8} + 4\frac{5}{8} = 4+\frac{7}{8} + 4+\frac{5}{8} =\frac{19}{2} \]